# A Construction of Infinite Families of Self-Orthogonal Quasi-Cyclic Codes Using Constituent Codes

### v1

@@ 2.3. A Lower Bound of the Minimum Distance of a QC Code
Let $C_{1}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{1}}}),...,C_{h}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{h}}})$ be the nonzero constituent codes of a $\ell$-QC code $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$. By [16, Lemma 4.1], the trace-formula described in (13) may be taken from an specific finite extension field $\mathbb{F}_{q^{w}}$, where $\mathbb{F}_{q}(\alpha^{u_{a_{i}}})\subset\mathbb{F}_{q^{w}}$, for all $1\leq i\leq h$, given the existence of elements $w_{i}\in\mathbb{F}_{q}(\alpha^{u_{a_{i}}})$ so that $\operatorname{Tr}_{\mathbb{F}_{q^{w}}/\mathbb{F}_{q}(\alpha^{u_{a_{i}}})}(w_{i})=1$. In this case, looking $\mathcal{C}$ as a collection of arrays, the columns of any codeword $\textbf{c}\in\mathcal{C}$ are codewords of a $m$-length cyclic code $D:=D_{1,2,..,h}$, whose dual’s basic zero set is
@@
In [39], the authors called $D$ the cyclic code associated to $\mathcal{C}$. Let $I=\{i_{1},i_{2},...,i_{t}\}$ be a nonempty subset of $\{1,2,...,h\}$ where
@@
The next result uses the $R_{I}$ to provide a lower bound to the minimum distance of a $\ell$-QC code $\mathcal{C}$ based on the minimum distances of its constituent codes and the cyclic code (and its cyclic subcodes) associated to $\mathcal{C}$. This lower bound is usually referred as GO-bound where relevant details can be checked in [16]. First, we assume that the minimum distances of the nonzero constituent codes $C_{1},C_{2},...,C_{h}$ of $\mathcal{C}$ satisfy the following condition
@@
[16, Theorem 4.8] Let $\mathcal{C}$ be a QC code as in (10) with the assumption (25). Then we have
@@ 2.4. Constructions of Quantum Error-Correcting Codes from Classical Ones
From [23], the use of classical codes has became one of the standard methods for constructing quantum error-correcting codes. A quantum error-correcting code (QECC) $Q$ is a $K$-dimensional subspace of $(\mathbb{C}^{q})^{\otimes n}$. If $Q$ has minimum distance $d$, then we say that $Q$ is an $n,K,d]]_{q}$ code. If $K=q^{k}$, we write $n,k,d]]_{q}$. The length $n$, the dimension $K$, and minimum distance $d$ are the parameters of $Q$.
If $Q$ is a quantum code with minimum distance $d$, $Q$ can detect up to d1d-1 quantum errors, and can correct up to $\lfloor\frac{d-1}{1}\rfloor$ quantum errors. The celebrated Singleton bound [24, Lemma 3.5.1] says that the parameters of an $n,K,d]]_{q}$ quantum code satisfy
@@
This paper focuses on the construction and existence of quantum stabilizer codes, where a stabilizer (quantum) code $Q\neq\{\textbf{0}\}$ is the common eigenspace of a commutative subgroup of the error group generated by a nice basis of $(\mathbb{C}^{q})^{\otimes n}$. The code $Q$ is said to be pure to d, or simply pure, if and only if its stabilizer group does not contain non-scalar matrices of weight less than its minimum distance $d$. (See Definitions 3.5.7 and 3.5.8 in [24].) Moreover, a code $Q$ is said to be impure if and only if there are non-identity stabilizer elements of weight less than the minimum distance. In 1996, the following construction was introduced by Calderbank and Shor [6], and Steane [38], and is notably the most direct link between classical and quantum coding theory.
Theorem 2.10 (CSS construction [23]).
@@
[2, 23] Let $C$ be a classical $[n,k,d]$ code over $\mathbb{F}_{q}$. If $C$ is dual-containing, then there exists a quantum stabilizer code $Q$ with parameters $n,2k-n,\geq d]]_{q}$ that is pure to $d$. Moreover, if the minimum distance of $C^{\perp}$ exceeds $d$, then the quantum code $Q$ has minimum distance $d$.
In recent years, researchers in coding theory have worked diligently to construct good (e.g., quantum MDS) quantum codes from classical Euclidean, Hermitian, or symplectic self-orthogonal, dual-containing, or self-dual codes [1, 3, 7, 10, 11, 12, 13, 14, 21, 27, 34, 40, 41]. Moreover, the following results (which will be used in examples in Section 4) present new quantum code constructions utilizing the lengthening, shortening, and dimension reduction of a given stabilizer quantum code.
@@
[23, Lemmas 69, 70, 71, and Corollary 73]
@@ 3.1.1. Recursive Construction:
for all $1\leq i\leq s-1$. Note that repeated application of (47) yields $d(C_{ui})\geq m^{u-1}d(C_{1i})$ for each $i$. This condition can be satisfied, for instance, by setting $C_{ui}=mC_{u-1,i}$. However, unlike in Lemma 3.1, taking copies of constituent codes is not required to obtain our result.
@@
where $d(D_{1,2,...,2t+s-1})-d(D_{1,2,...,2t+s})>0$ from Equation (4.19) in [16]. Therefore by Theorem 2.9,
@@
where $\Phi_{d}$ is the $d$-th cyclotomic polynomial, there would be at least four cyclotomic polynomial factors ($1$, $m$, and $m$’s distinct prime factors). This would result in at least four irreducible factors. If $m$ is prime,
@@ 3.2. Constructing ESO/ESD/EDC QC Codes with Square-Root-Like Lower Bound
Self-dual codes have applications in several mathematical problems such as in the construction of designs, lattices, and quantum error-correcting codes (see Subsection 2.4). Notably, the construction of infinite families of $n_{i},k_{i},d_{i}]_{q}$ self-dual codes is itself an interesting problem. Additionally, code families where $d_{i}\geq c\sqrt{n_{i}}$ for some constant $c$ are said to satisfy the square-root-like lower bound. In the special case when $c=1$, we say that the code family satisfies the square-root lower bound. According to [8], there are very few constructions of infinite families of self-dual codes whose minimum distances meet the square-root lower bound, which are depicted in that paper. Notably, the work in  [5] presents a unique construction of an infinite family of self-dual QC codes.
To the best of our knowledge, the authors of [8] in 2025 presented the first infinite families of non-binary ESD (and HSD) cyclic codes satisfying the square-root-like lower bound to their minimum distances. Using results from [26] and developments presented in this work (see Corollary 3.4), we construct several infinite families of ESD QC codes which satisfy the square-root like lower bound. Moreover, we also construct HSD QC codes using results in the next subsection. We first state the following helpful lemma.
@@
We conclude this subsection by emphasizing that, alongside the seminal constructions of infinite families of self-dual cyclic codes satisfying the square-root-like bound in [8], our work introduces a novel construction of infinite families of QC codes that also attain this square-root-like lower bound on minimum distance.
3.3. Galois Closed QC Codes and Their Duals
Having QC codes whose Hermitian and Euclidean duals are equal is a valuable property as it allows us to move freely between the two inner products. To achieve this property, we introduce the notion of the Galois closure of a code $C$: By  [20], given a positive integer $m\geq 1$ and $c=(c_{1},c_{2},...,c_{n})\in C\subset\mathbb{F}_{q}^{n}$ (in this case, $C$ is just a subset), we define $c^{m}=(c_{1}^{m},c_{2}^{m},...,c_{n}^{m})$ and $C^{m}=\{c^{m}:c\in C\}\subset\mathbb{F}_{q}^{n}.$ If $C$ is a subspace of $\mathbb{F}_{q}^{n}$ and $m$ is a power of the characteristic of $\mathbb{F}_{q}$, i.e., $m=q^{t}$ for some $t\in\mathbb{Z}^{+}$, then $C^{m}$ is also a subspace of $\mathbb{F}_{q}^{n}$.
Let $\mathbb{F}_{q}\subset\mathbb{F}_{q^{r}}$ and $C$ be an $\mathbb{F}_{q^{r}}$-code. We call $C$ Galois closed if $C=C^{q}$. See [4] for further information. The Frobenius map $C\mapsto C^{q}$ respects the quasi-cyclic property of $C$ so that $C^{q}$ is also quasi-cyclic whenever $C$ is. Furthermore, if $\phi(C),\phi(C^{q})\subset\left(\frac{\mathbb{F}_{q^{2}}[x]}{\langle x^{m}-1\rangle}\right)^{\ell}$, then we can look at the constituent codes of $C$ and $C^{q}$ and assume that the irreducible factors of $x^{m}-1$ are aligned for both decompositions.
@@ 3.3. Galois Closed QC Codes and Their Duals
Let $C\subseteq\mathbb{F}_{q^{2}}^{\ell}$ be a $k$-dimensional code. Then $C$ is Galois closed if and only if $C=\bigoplus_{i}^{k}C_{i}$ where $C_{i}=C_{i}^{q}$ and $\dim(C_{i})=1$, i.e., $C$ is Galois closed if any only if can be decomposed as Galois closed codes of dimension $1$.
@@ 4. Constructions of QECCs from QC Codes
Given the construction of $\ell$-QC codes and results seen in [16, 26], along with the collection of quantum constructions in Subsection 2.4, we present in this section a construction of quantum codes with good parameters.
@@
Let $q=4$, $\ell=3$, $m=7$, and $\alpha$ be a primitive $7$-th root of unity. Let
@@
For any element $\gamma\in\mathbb{F}_{4}(\alpha)\setminus\mathbb{F}_{4}$, let $C_{11}^{\prime}=\left\langle(\gamma,\gamma,\gamma)\right\rangle\subset\mathbb{F}_{4}^{3}(\alpha)$ , $C_{11}^{\prime\prime}=(C_{11}^{\prime})^{\perp}\subset\mathbb{F}_{4}^{3}(\alpha^{3})$, and $C_{11}=\mathbb{F}_{4}^{3}$ be the constituent codes, whose parameters are $3,1,3]_{64}$, $3,2,2]_{64}$, and $3,3,1]_{4}$, respectively. Notice that $C_{11}^{\prime},C_{11}^{\prime\prime}$, and $C_{11}$ are Galois closed by Proposition 3.12, and $C_{11}$ is trivially HDC. Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective $3$-QC-code which by Theorem 3.13 is also Galois-closed. The minimum distance of the cyclic code associated to $\mathcal{C}$ and all its associated cyclic subcodes are described below
@@
and $\mathcal{C}$ is a $21,12,\geq 7]_{4}-$EDC $\ell$-QC code by [26, Theorem 4.2]. As seen in the Corollary 2.11, from $\mathcal{C}$ it is possible to obtain a $21,3,d^{\perp}\geq 7]]_{4}-$quantum code $Q$. According to [15], the best known quaternary QECC for $n=21$ and $k=3$ has minimum distance $7.$
@@
Code Table [15] 6–7 6–8 7–8 7–8 8–9
While there is no marked improvement on the minimum distance for all of the above parameters, we see that this explicit code is at least close to the best minimum distance for most. We do, however, obtain improved lower bounds for $[[22,3]]$ and $[[23,3]]$. Indeed in the first case, we have that 7 is attainable as a minimum distance, and in the second case, the best minimum distance is now 7–8 rather than 6–8.
@@
Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective $3$-QC-code. By Theorem 3.7, we have an infinity family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of EDC QC codes whose their parameters are
@@
Let $q=2$, $m=7$, $\ell=8$, and $\alpha$ be a primitive $7$-th root of unity. Assume the following decomposition of $R_{2,7}^{8}=\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{7}-1\rangle}\right)^{8}$. Thus
@@
Let $C_{11}^{\prime}$ a $8,3,6]_{8}$ MDS code (its construction can be checked in [15]), $C_{11}^{{}^{\prime\prime}}=C_{11}^{{}^{\prime}\perp}$ a $8,5,4]_{8}$ (also MDS) code, and $C_{11}$ be a EDC $8,6,2]_{2}-$code with basis
@@
From the minimum distances of the associated cyclic subcodes provided in the Example 4.1, together with the minimum distances of the constituent codes described above, the minimum distance of $\mathcal{C}$ in $\mathbb{F}_{2}^{56}$ is $d\geq d_{GO}=14.$ Moreover, by Lemma 2.1, the dimension of $\mathcal{C}$ is 30, namely, $\mathcal{C}$ is a $8$-QC EDC $56,30,\geq 14]_{2}-$code and, based on Corollary 2.11, it is possible to construct a QECC with parameters $56,4,\geq 14]]_{2}$. Notice that in [15], the best $56,4]]_{2}-$QECC has minimum distance 13.
By the Lemma 2.12, we list more good quantum codes from this code. Similar to the above, we have an improvement on some lower bounds from this explicit example: the best minimum distance for $[[55,5]]$ can be improved to 13–18 and $[[57,4]]$ can be improved to 14–18.
$[[51,9]]$ $[[52,8]]$ $[[53,7]]$ $[[54,6]]$ $[[55,5]]$

### v2

@@ 2.3. A Lower Bound of the Minimum Distance of a QC Code
Let $C_{1}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{1}}}),...,C_{h}\subset\mathbb{F}_{q}^{\ell}(\alpha^{u_{a_{h}}})$ be the nonzero constituent codes of a $\ell$-QC code $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$. By [cem, Lemma 4.1], the trace-formula described in (13) may be taken from an specific finite extension field $\mathbb{F}_{q^{w}}$, where $\mathbb{F}_{q}(\alpha^{u_{a_{i}}})\subset\mathbb{F}_{q^{w}}$, for all $1\leq i\leq h$, given the existence of elements $w_{i}\in\mathbb{F}_{q}(\alpha^{u_{a_{i}}})$ so that $\operatorname{Tr}_{\mathbb{F}_{q^{w}}/\mathbb{F}_{q}(\alpha^{u_{a_{i}}})}(w_{i})=1$. In this case, looking $\mathcal{C}$ as a collection of arrays, the columns of any codeword $\textbf{c}\in\mathcal{C}$ are codewords of a $m$-length cyclic code $D:=D_{1,2,..,h}$, whose dual’s basic zero set is
@@
In [BastosAlvarezFloresSalerno], the authors called $D$ the cyclic code associated to $\mathcal{C}$. Let $I=\{i_{1},i_{2},...,i_{t}\}$ be a nonempty subset of $\{1,2,...,h\}$ where
@@
The next result uses the $R_{I}$ to provide a lower bound to the minimum distance of a $\ell$-QC code $\mathcal{C}$ based on the minimum distances of its constituent codes and the cyclic code (and its cyclic subcodes) associated to $\mathcal{C}$. This lower bound is usually referred as GO-bound where relevant details can be checked in [cem]. First, we assume that the minimum distances of the nonzero constituent codes $C_{1},C_{2},...,C_{h}$ of $\mathcal{C}$ satisfy the following condition
@@
[cem, Theorem 4.8] Let $\mathcal{C}$ be a QC code as in (10) with the assumption (25). Then we have
@@ 2.4. Constructions of Quantum Error-Correcting Codes from Classical Ones
From [Ketkar-Klappenecker-Kumar-Sarvepalli2006], the use of classical codes has became one of the standard methods for constructing quantum error-correcting codes. A quantum error-correcting code (QECC) $Q$ is a $K$-dimensional subspace of $(\mathbb{C}^{q})^{\otimes n}$. If $Q$ has minimum distance $d$, then we say that $Q$ is an $n,K,d]]_{q}$ code. If $K=q^{k}$, we write $n,k,d]]_{q}$. The length $n$, the dimension $K$, and minimum distance $d$ are the parameters of $Q$.
If $Q$ is a quantum code with minimum distance $d$, $Q$ can detect up to d1d-1 quantum errors, and can correct up to $\lfloor\frac{d-1}{1}\rfloor$ quantum errors. The celebrated Singleton bound [LaGuardia2021, Lemma 3.5.1] says that the parameters of an $n,K,d]]_{q}$ quantum code satisfy
@@
This paper focuses on the construction and existence of quantum stabilizer codes, where a stabilizer (quantum) code $Q\neq\{\textbf{0}\}$ is the common eigenspace of a commutative subgroup of the error group generated by a nice basis of $(\mathbb{C}^{q})^{\otimes n}$. The code $Q$ is said to be pure to d, or simply pure, if and only if its stabilizer group does not contain non-scalar matrices of weight less than its minimum distance $d$. (See Definitions 3.5.7 and 3.5.8 in [LaGuardia2021].) Moreover, a code $Q$ is said to be impure if and only if there are non-identity stabilizer elements of weight less than the minimum distance. In 1996, the following construction was introduced by Calderbank and Shor [Calderbank-Shor1996], and Steane [Steane1996], and is notably the most direct link between classical and quantum coding theory.
Theorem 2.10 (CSS construction [Ketkar-Klappenecker-Kumar-Sarvepalli2006]).
@@
[Aly-Klappenecker-Sarvepalli2007, Ketkar-Klappenecker-Kumar-Sarvepalli2006] Let $C$ be a classical $[n,k,d]$ code over $\mathbb{F}_{q}$. If $C$ is dual-containing, then there exists a quantum stabilizer code $Q$ with parameters $n,2k-n,\geq d]]_{q}$ that is pure to $d$. Moreover, if the minimum distance of $C^{\perp}$ exceeds $d$, then the quantum code $Q$ has minimum distance $d$.
In recent years, researchers in coding theory have worked diligently to construct good (e.g., quantum MDS) quantum codes from classical Euclidean, Hermitian, or symplectic self-orthogonal, dual-containing, or self-dual codes [Abdukhalikov-Bag-Panario2023, Ball21, Cao24, DinhLeNguyenTansuchat21, FangFu19, GalindoHernando, galindo, GaoSunYanZhao25, JinXing13, LiuLiu21, ShiYueChang18, TianZhu19, WanZhengZhu24]. Moreover, the following results (which will be used in examples in Section 4) present new quantum code constructions utilizing the lengthening, shortening, and dimension reduction of a given stabilizer quantum code.
@@
[Ketkar-Klappenecker-Kumar-Sarvepalli2006, Lemmas 69, 70, 71, and Corollary 73]
@@ 3.1.1. Recursive Construction:
for all $1\leq i\leq s-1$. Note that repeated application of (47) yields $d(C_{ui})\geq m^{u-1}d(C_{1i})$ for each ii. This condition can be satisfied, for instance, by setting $C_{ui}=mC_{u-1,i}$. However, unlike in Lemma 3.1, taking copies of constituent codes is not required to obtain our result.
@@
where $d(D_{1,2,...,2t+s-1})-d(D_{1,2,...,2t+s})>0$ from Equation (4.19) in [cem]. Therefore by Theorem 2.9,
@@
where $\Phi_{d}$ is the $d$-th cyclotomic polynomial, there would be at least four cyclotomic polynomial factors (11, $m$, and $m$’s distinct prime factors). This would result in at least four irreducible factors. If $m$ is prime,
@@ 3.2. Constructing ESO/ESD/EDC QC Codes with Square-Root-Like Lower Bound
Self-dual codes have applications in several mathematical problems such as in the construction of designs, lattices, and quantum error-correcting codes (see Subsection 2.4). Notably, the construction of infinite families of $n_{i},k_{i},d_{i}]_{q}$ self-dual codes is itself an interesting problem. Additionally, code families where $d_{i}\geq c\sqrt{n_{i}}$ for some constant $c$ are said to satisfy the square-root-like lower bound. In the special case when $c=1$, we say that the code family satisfies the square-root lower bound. According to [infiniteselfdualfamilies], there are very few constructions of infinite families of self-dual codes whose minimum distances meet the square-root lower bound, which are depicted in that paper. Notably, the work in  [Calderbank] presents a unique construction of an infinite family of self-dual QC codes.
To the best of our knowledge, the authors of [infiniteselfdualfamilies] in 2025 presented the first infinite families of non-binary ESD (and HSD) cyclic codes satisfying the square-root-like lower bound to their minimum distances. Using results from [lingsole] and developments presented in this work (see Corollary 3.4), we construct several infinite families of ESD QC codes which satisfy the square-root like lower bound. Moreover, we also construct HSD QC codes using results in the next subsection. We first state the following helpful lemma.
@@
We conclude this subsection by emphasizing that, alongside the seminal constructions of infinite families of self-dual cyclic codes satisfying the square-root-like bound in [infiniteselfdualfamilies], our work introduces a novel construction of infinite families of QC codes that also attain this square-root-like lower bound on minimum distance.
$3.$ Galois Closed QC Codes and Their Duals
Having QC codes whose Hermitian and Euclidean duals are equal is a valuable property as it allows us to move freely between the two inner products. To achieve this property, we introduce the notion of the Galois closure of a code $C$: By  [Jin2010], given a positive integer $m\geq 1$ and $c=(c_{1},c_{2},...,c_{n})\in C\subset\mathbb{F}_{q}^{n}$ (in this case, $C$ is just a subset), we define $c^{m}=(c_{1}^{m},c_{2}^{m},...,c_{n}^{m})$ and $C^{m}=\{c^{m}:c\in C\}\subset\mathbb{F}_{q}^{n}.$ If $C$ is a subspace of $\mathbb{F}_{q}^{n}$ and $m$ is a power of the characteristic of $\mathbb{F}_{q}$, i.e., $m=q^{t}$ for some $t\in\mathbb{Z}^{+}$, then $C^{m}$ is also a subspace of $\mathbb{F}_{q}^{n}$.
Let $\mathbb{F}_{q}\subset\mathbb{F}_{q^{r}}$ and $C$ be an $\mathbb{F}_{q^{r}}$-code. We call $C$ Galois closed if $C=C^{q}$. See [Bierbrauer] for further information. The Frobenius map $C\mapsto C^{q}$ respects the quasi-cyclic property of $C$ so that $C^{q}$ is also quasi-cyclic whenever $C$ is. Furthermore, if $\phi(C),\phi(C^{q})\subset\left(\frac{\mathbb{F}_{q^{2}}[x]}{\langle x^{m}-1\rangle}\right)^{\ell}$, then we can look at the constituent codes of $C$ and $C^{q}$ and assume that the irreducible factors of $x^{m}-1$ are aligned for both decompositions.
@@
Let $C\subseteq\mathbb{F}_{q^{2}}^{\ell}$ be a $k$-dimensional code. Then $C$ is Galois closed if and only if $C=\bigoplus_{i}^{k}C_{i}$ where $C_{i}=C_{i}^{q}$ and $\dim(C_{i})=1$, i.e., $C$ is Galois closed if any only if can be decomposed as Galois closed codes of dimension 11.
@@ 4. Constructions of QECCs from QC Codes
Given the construction of $\ell$-QC codes and results seen in [cem, lingsole], along with the collection of quantum constructions in Subsection 2.4, we present in this section a construction of quantum codes with good parameters.
@@
Let $q=4$, $\ell=3$, $m=7$, and $\alpha$ be a primitive 77-th root of unity. Let
@@
For any element $\gamma\in\mathbb{F}_{4}(\alpha)\setminus\mathbb{F}_{4}$, let $C_{11}^{\prime}=\left\langle(\gamma,\gamma,\gamma)\right\rangle\subset\mathbb{F}_{4}^{3}(\alpha)$ , $C_{11}^{\prime\prime}=(C_{11}^{\prime})^{\perp}\subset\mathbb{F}_{4}^{3}(\alpha^{3})$, and $C_{11}=\mathbb{F}_{4}^{3}$ be the constituent codes, whose parameters are $3,1,3]_{64}$, $3,2,2]_{64}$, and $3,3,1]_{4}$, respectively. Notice that $C_{11}^{\prime},C_{11}^{\prime\prime}$, and $C_{11}$ are Galois closed by Proposition 3.12, and $C_{11}$ is trivially HDC. Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective 33-QC-code which by Theorem 3.13 is also Galois-closed. The minimum distance of the cyclic code associated to $\mathcal{C}$ and all its associated cyclic subcodes are described below
@@
and $\mathcal{C}$ is a $21,12,\geq 7]_{4}-$EDC $\ell$-QC code by [lingsole, Theorem 4.2]. As seen in Corollary 2.11, from $\mathcal{C}$ it is possible to obtain a $21,3,d^{\perp}\geq 7]]_{4}-$quantum code $Q$. According to [table], the best known quaternary QECC for $n=21$ and $k=3$ has minimum distance $7.$
@@
Code Table [table] 7-10 8-11 8-11 9-11 9-12
Although there is no significant improvement in the minimum distances for any of the above parameters, most of our codes achieve values that are within one or two of the best-known lower bounds. It is also worth noting that our example exhibits lower bounds of the minimum distance, so it may be possible to obtain codes with larger minimum distances.
@@
Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective 33-QC-code. By Theorem 3.7, we have an infinity family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of EDC QC codes whose their parameters are
@@
Let $q=2$, $m=7$, $\ell=8$, and $\alpha$ be a primitive 77-th root of unity. Assume the following decomposition of $R_{2,7}^{8}=\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{7}-1\rangle}\right)^{8}$. Thus
@@
Let $C_{11}^{\prime}$ be a $8,3,6]_{8}$ MDS code (its construction can be checked in [table]), $C_{11}^{{}^{\prime\prime}}=C_{11}^{{}^{\prime}\perp}$ a $8,5,4]_{8}$ (also MDS) code, and $C_{11}$ be a EDC $8,6,2]_{2}-$code with basis
@@
From the minimum distances of the associated cyclic subcodes provided in the Example 4.1, together with the minimum distances of the constituent codes described above, the minimum distance of $\mathcal{C}$ in $\mathbb{F}_{2}^{56}$ is $d\geq d_{GO}=14.$ Moreover, by Lemma 2.1, the dimension of $\mathcal{C}$ is 30, namely, $\mathcal{C}$ is a 88-QC EDC $56,30,\geq 14]_{2}-$code and, based on Corollary 2.11, it is possible to construct a QECC with parameters $56,4,\geq 14]]_{2}$. Notice that in [table], the best $56,4]]_{2}-$QECC has minimum distance 13.
By the Lemma 2.12, we list more good quantum codes from this code. We have an improvement on some lower bounds from this explicit example: the best minimum distance for $55,5]]_{2}$ can be improved to 13–18 and $57,4]]_{2}$ can be improved to 14–19.
$51,9]]_{2} [[$$52,8]]_{2} [[$$53,7]]_{2} [[$$4,6]]_{2} [[55$$5,5]]_{2}$